A bag contains five balls numbered 1,2,3,4,5. Another bag contains six balls numbered 1,2,3,4,5,6. One ball is drawn at random from each bag. Find the probability that (i) one ball is numbered 1 and the other 6, (ii) both balls have an odd number.
Answer by stanbon(75887)

A bag contains five balls numbered 1,2,3,4,5.
Another bag contains six balls numbered 1,2,3,4,5,6.
One ball is drawn at random from each bag.
Find the probability that
(i) P(one ball is numbered 1 and the other 6) = (1/5)(1/6) = 1/30
-------------------
(ii) P(both balls have an odd number) = (3/5)(3/6) = 9/30 = 3/10
======================
Cheers,
Stan H.
======================